OpenStudy (anonymous):

Find the values of the trigonometric functions of t from the given information. sec(t) = 6, sin(t) < 0 need sin(t, cos(t), csc(t), tan(t), cot(t)

OpenStudy (anonymous):

\[\sec(t) = 1/\cos(t)\] so:\[1/\cos(t) = 6 \]\[\cos(t) = 1/6\] then: \[c^2 = x^2 + y^2\] \[y^2 = c^2 - x^2 = (36)-1 = \sqrt{35}\] so: \[\sin(t) = \sqrt{35}\] \[\csc(t) = 1/\sin(t) = 1/\sqrt{35}\] \[\tan(t) = \sin(t)/\cos(t) =6\sqrt{35}\] \[\cot(t) = 1/\tan(t) = 1/6\sqrt{35}\] ^_^

OpenStudy (anonymous):

wait the hyp =36 so : sin(t) = sqrt(35)/36 csc(t) = 36/sqrt(35) tan(t) = sqrt(35)/6 cot(t) = 6/sqrt(35) Sorry for the mess ^^" .